# Coulomb Branch Operator Algebras and Universal Selection Rules for $\mathcal{N}=2$ SCFTs

### v1

@@ 5 Discussion
    $\mathcal{A}_{\mathcal{C}S}:={\mathcal{A}_{\mathcal{C}}\times\mathcal{A}_{\rm Schur}\over\mathcal{A}_{\mathcal{C}}\cap\mathcal{A}_{\rm Schur}}~{},$ (5.1)
    where $\mathcal{A}_{\mathcal{C}}\cap\mathcal{A}_{\rm Schur}$ is the non-trivial sub-algebra common to the Coulomb branch and Schur operator algebras (locality implies that $\hat{\mathcal{C}}_{0(0,0)}\in\mathcal{A}_{\mathcal{C}}\cap\mathcal{A}_{\rm Schur}$, and so the corresponding subalgebra generated by the EM tensor multiplet satisfies $\langle\hat{\mathcal{C}}_{0(0,0)}\rangle\leq\mathcal{A}_{\mathcal{C}}\cap\mathcal{A}_{\rm Schur}$). In particular, it would be interesting to understand when this combined algebra saturates the local operator algebra of the SCFT, $\mathcal{A}_{\mathcal{C}S}=\mathcal{A}_{\mathcal{O}}$ (i.e., when the Coulomb branch and Schur sector operators generate the local operator algebra). [^30] Note that in the case of Coulombic theories, $\mathcal{A}_{\rm Schur}<\mathcal{A}_{\mathcal{C}}$ and so $\mathcal{A}_{\mathcal{C}S}=\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$. Note that this algebra is sensitive to the continuous flavor symmetries of 4d $\mathcal{N}=2$ SCFTs (unlike $\mathcal{A}_{\mathcal{C}}$), although, a priori, it is logically possible that it may sometimes fail to capture the global form of the continuous flavor symmetry of $\mathcal{A}_{\mathcal{O}}$.
    As a final note regarding (5.1), we observe that it is natural to relate $\mathcal{A}_{\mathcal{C}S}$ to the structure of moduli spaces, and it would be intriguing to understand if this algebra can be used to give a more intrinsic SCFT definition of the moduli space.

### v2

@@ 5 Discussion
    $\mathcal{A}_{\mathcal{C}S}:=\mathcal{A}_{\mathcal{C}}\times\mathcal{A}_{\rm Schur}~{}.$ (5.1)
    What we mean by the above expression is the operator algebra generated by the Coulomb branch multiplets and the Schur multiplets. [^31] Note that in the case of Coulombic theories, $\mathcal{A}_{\rm Schur}<\mathcal{A}_{\mathcal{C}}$ and so $\mathcal{A}_{\mathcal{C}S}=\mathcal{A}_{\mathcal{C}}$. More generally, this statement need not hold. This combined operator algebra characterizes many important properties of 4d $\mathcal{N}=2$ SCFTs. [^32] Note that Coulomb branch operators and Schur operators, while describing different physics, “know” a lot about each other (for some initial results in this direction, see [44, 51]; more recently see also [52]). It would be interesting to understand this relationship in full generality. In particular, it would be interesting to understand when this combined algebra saturates the local operator algebra of the SCFT, $\mathcal{A}_{\mathcal{C}S}=\mathcal{A}_{\mathcal{O}}$ (i.e., when the Coulomb branch and Schur sector operators generate the local operator algebra). [^33] In Coulombic theories, $\mathcal{A}_{\rm Schur}<\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$ and so $\mathcal{A}_{\mathcal{C}S}=\mathcal{A}_{\mathcal{O}}$. Note that this algebra is sensitive to the continuous flavor symmetries of 4d $\mathcal{N}=2$ SCFTs (unlike $\mathcal{A}_{\mathcal{C}}$), although, a priori, it is logically possible that it may sometimes fail to capture the global form of the continuous flavor symmetry of $\mathcal{A}_{\mathcal{O}}$.
    As a final note regarding (5.1), we observe that it is natural to relate $\mathcal{A}_{\mathcal{C}S}$ to the structure of moduli spaces, and it would be intriguing to understand if this algebra can be used to give a more intrinsic SCFT definition of the moduli space. Note that the Schur sector naturally captures properties of the Higgs branch, and so we expect all moduli spaces of 4d $\mathcal{N}=2$ SCFTs are described by operators in (5.1).